# Discovery potential of a long-lived partner of inelastic dark matter at MATHUSLA in $U(1)_{X_{3}}$ extension of the standard model

### v1

@@ IV Prospect for MATHUSLA
The travel distance $c\tau\simeq 100$ m is the most sensitive range at MATHUSLA  Chou:2016lxi ; Curtin:2017izq ; Curtin:2017bxr . The pair production cross section of such a LLP $\chi$ for its $5\sigma$ discovery can be read as $\sigma_{\chi\chi}>0.3$ fb from Fig. 1 in Ref. Jana:2018rdf . On the other hand, in our scenario, the LLP $P$ is not pair produced but singly produced through $pp\rightarrow Z^{\prime}\rightarrow SP$. Thus, the required production cross section for the LLP $P$ to be discovered at the MATHUSLA would be set to be larger than $0.6$ fb. In Tab. 3, we list the $SP$ production cross section at $13$ TeV LHC and $14$ TeV LHC for our benchmark on Tab. 2 satisfying the DM abundance and the LHC bounds. We find that the MATHUSLA will be able to discover those LLPs because the production cross section $\sigma_{SP}$ is sufficiently larger than $0.6$ fb.
@@
$BP_{-2}$ $154$ fb ($<203$ fb) $10.8$ fb $12.9$ fb
Table 3:
@@ V Summary
We have proposed a simple extension of the SM with an extra $U(1)_{R_{3}}$ gauge interaction with scalar particle of DM candidate with the charge $1$, third generation of SM fermions and third generation of right-handed neutrinos. After the SM singlet $U(1)_{R_{3}}$ breaking scalar with the charge $2$ develops the VEV, the gauge boson acquires the mass, and the tiny mass splitting between real and imaginary component of the charge $1$ scalar appears through the scalar tri-linear interaction. The lighter scalar $S$ is inelastic DM candidate with the heavier state $P$. Due to the mass degeneracy, the slightly heavier state $P$ is long-lived and will be able to be discovered as LLPs.
We have calculated the production cross section of LLP $P$ through the process $pp\rightarrow Z^{\prime}\rightarrow SP$ at the MATHUSLA for a benchmark point which satisfies stringent LHC bounds and thermal DM abundance. The production cross section of benchmark point turns out to be about $10$ fb, which is about an order of magnitude larger than the cross section required by $5\sigma$ discovery as a LLP at the MATHUSLA. Thus, we conclude that the heavier state of inelastic DM in our model can be discovered at the MATHUSLA.
@@ Acknowledgments
Appendix A The decay rate of $P$

### v2

@@ IV Prospect for MATHUSLA
The travel distance $c\tau\simeq 100$ m is the most sensitive range at MATHUSLA  Chou:2016lxi ; Curtin:2017izq ; Curtin:2017bxr . The pair production cross section of such an LLP $\chi$ for its $5\sigma$ discovery can be read as $\sigma_{\chi\chi}>0.3$ fb from Fig. 1 in Ref. Jana:2018rdf . On the other hand, in our scenario, the LLP $P$ is not pair produced but singly produced through $pp\rightarrow Z^{\prime}\rightarrow SP$. Thus, the required production cross section for the LLP $P$ to be discovered at MATHUSLA would be set to be larger than $0.6$ fb. In Table 3, we list the $SP$ production cross section at $13$ TeV LHC and $14$ TeV LHC for our benchmark on Table 2 satisfying the DM abundance and the LHC bounds. We find that MATHUSLA will be able to discover those LLPs because the production cross section $\sigma_{SP}$ is sufficiently larger than $0.6$ fb.
@@
$BP_{1}$ $154$ fb ($<203$ fb) $10.8$ fb $12.9$ fb
$BP_{2}$ $7.78$ fb ($<8.24$ fb) $0.697$ fb $0.888$ fb
Table 3: Estimated production cross section for tau pairs and $SP$ pairs in benchmark points.
@@ V Summary
We have proposed a simple extension of the SM with an extra $U(1)_{R_{3}}$ gauge interaction with a scalar particle of DM candidate with the charge $1$, third generation of SM fermions and third generation of right-handed neutrinos. After the SM singlet $U(1)_{R_{3}}$ breaking scalar with charge $2$ develops a VEV, the gauge boson acquires a mass, and a tiny mass splitting between the real and imaginary components of the charge $1$ scalar appears through the scalar trilinear interaction. The lighter scalar $S$ is an inelastic DM candidate with a heavier state $P$. Due to the mass degeneracy, the slightly heavier state $P$ is long-lived and will be able to be discovered as LLPs.
We have calculated the production cross section of LLP $P$ through the process $pp\rightarrow Z^{\prime}\rightarrow SP$ at MATHUSLA for a benchmark point that satisfies stringent LHC bounds and thermal DM abundance. The production cross section of the benchmark point turns out to be about $10$ fb, which is about an order of magnitude larger than the cross section required by $5\sigma$ discovery as a LLP at the MATHUSLA. Thus, we conclude that the heavier state of inelastic DM in our model can be discovered at MATHUSLA.
@@ Acknowledgments
Appendix A An extension of Higgs sector
We note a possible UV completed model to generate top qurak mass and how gauge kinetic mixing parameter affects the mixing between $Z$ and $Z^{\prime}$. We may consider an extension with three Higgs doublet fields $\Phi_{1},\Phi_{2}$ and $\Phi_{3}$ and some singlet scalar fields to eliminate unwanted NG modes. The charge assignments of those Higgs fields are in Tab. 4. Those doublet fields develop the VEVs as
$\Phi_{i}=$ $\left(\begin{array}[]{c}0\\ \frac{v_{i}}{\sqrt{2}}\\ \end{array}\right).$ (26)
The quark Yukawa interactions are expressed as
$\mathcal{L}_{\mathrm{Yukawa}}=$ $\sum_{i,j}^{1,2}\left(-y_{ij}^{u}\overline{Q}_{i}\tilde{\Phi}_{1}u_{jR}-y_{ij}^{d}\overline{Q}_{i}\Phi_{1}d_{jR}-y_{3j}^{u}\overline{Q}_{3}\tilde{\Phi}_{1}u_{jR}-y_{3j}^{d}\overline{Q}_{3}\Phi_{1}d_{jR}\right)$
$-y_{i3}^{u}\overline{Q}_{i}\tilde{\Phi}_{2}u_{3R}-y_{i3}^{d}\overline{Q}_{i}\Phi_{2}d_{3R}-y_{33}^{u}\overline{Q}_{3}\tilde{\Phi}_{2}u_{3R}-y_{33}^{d}\overline{Q}_{3}\Phi_{2}d_{3R}+\mathrm{H.c.},$ (27)
and there are non-vanishing entries for all $3$ by $3$ elements. The mass of top quark is generated through the VEV $v_{2}$. At the same time, the mass matrix of neutral gauge bosons $\hat{Z}$ and $\hat{X}$ in the interaction basis are given by
$\left(\begin{array}[]{cc}M_{\hat{Z}\hat{Z}}^{2}&M_{\hat{Z}\hat{X}}^{2}\\ M_{\hat{X}\hat{Z}}^{2}&M_{\hat{X}\hat{X}}^{2}\\ \end{array}\right)=\left(\begin{array}[]{cc}\frac{g_{2}^{2}v^{2}}{8c_{W}^{2}}&-\frac{g_{2}g_{X}}{4c_{W}}\left(-v_{2}^{2}+q_{X_{3}}v_{3}^{2}\right)\\ -\frac{g_{2}g_{X}}{4c_{W}}\left(-v_{2}^{2}+q_{X_{3}}v_{3}^{2}\right)&\frac{g_{X}^{2}}{2}\left(v_{2}^{2}+q_{X_{3}}^{2}v_{3}^{2}+4v_{\phi}^{2}\right)\\ \end{array}\right),$ (28)
with $v^{2}=v_{1}^{2}+v_{2}^{2}+v_{3}^{2}$, where $g_{2}$ is the $SU(2)_{L}$ gauge coupling, $g_{X}$ is the $U(1)_{R_{3}}$ gauge coupling, $v_{\phi}$ is the VEV of singlet Higgs field with the $U(1)_{R_{3}}$ charge $-2$, the the $U(1)_{R_{3}}$ charges of $\Phi_{1},\Phi_{2}$ and $\Phi_{3}$ are taken to be $0,-1,q_{X_{3}}$, respectively and $c_{W}$ is the cosine of the Weinberg angle. Note that the mass mixing between $\hat{Z}$ and $\hat{X}$ is tiny for $-v_{2}^{2}+q_{X_{3}}v_{3}^{2}\simeq 0$ Seto:2020jal . The field redefinition by the orthogonal matrix resolve the gauge kinetic mixing
$\mathcal{L}_{\mathrm{Gauge}}\supset\frac{\sin\epsilon}{2}\hat{B}_{\mu\nu}\hat{X}^{\mu\nu},$ (29)
but induces additional mass mixings. The resultant mass matrix for the canonical kinetic terms becomes
$\left(\begin{array}[]{cc}M^{2}_{\hat{Z}\hat{Z}}&\frac{M^{2}_{\hat{Z}\hat{X}}}{c_{\epsilon}}-M^{2}_{\hat{Z}\hat{Z}}s_{W}t_{\epsilon}\\ \frac{M^{2}_{\hat{Z}\hat{X}}}{c_{\epsilon}}-M^{2}_{\hat{Z}\hat{Z}}s_{W}t_{\epsilon}&\frac{s_{W}c_{\epsilon}t_{\epsilon}}{c_{\epsilon}^{2}}\left(M^{2}_{\hat{Z}\hat{Z}}s_{W}c_{\epsilon}t_{\epsilon}-2M^{2}_{\hat{Z}\hat{X}}\right)+\frac{1}{c_{\epsilon}^{2}}M^{2}_{\hat{X}\hat{X}}\\ \end{array}\right).$ (30)
The magnitude of the off-diagonal component, or equvalently the $Z-Z^{\prime}$ mixing, is by experiments stringently constrained as $M^{2}_{\hat{Z}\hat{X}}-M^{2}_{\hat{Z}\hat{Z}}s_{W}s_{\epsilon}\lesssim(0.1\,\mathrm{GeV})^{2}$ ParticleDataGroup:2024cfk , so that the shift of the $Z$ boson mass due to the mixing must be within the error of its measurement. In our model, by seting the ratio $v_{2}/v_{3}$ in $M^{2}_{\hat{Z}\hat{X}}$ appropriately so that the off-diagonal element can be as small as to be consistent, the mixing between $Z$ and $Z^{\prime}$ is always negligible even under non-vanishing gauge kinetic parameter $\epsilon$.
One can easily see that physics of dark matter and $Z^{\prime}$ described in the main text are intact for the above extension.
SU(3)c SU(2)L U(1)Y U(1)${}_{R_{3}}$ $Z_{2}$
$Q^{i}$ 3 2 $1/6$ 00 $+$
$u^{i}_{R}$ 3 1 $2/3$ $-\delta_{i3}$ $+$
$d^{i}_{R}$ 3 1 $-1/3$ $\delta_{i3}$ $+$
$\Phi_{1}$ 1 2 $1/2$ 00 $+$
$\Phi_{2}$ 1 2 $1/2$ $-1$ $+$
$\Phi_{3}$ 1 2 $1/2$ $q_{X3}$ $+$
$\phi_{1}$ 1 1 00 $+1$-
$\phi_{1}^{\prime}$ 1 1 00 $+1$ $+$
$\phi_{2}$ 1 1 00 $+2$ $+$
$\phi_{3}$ 1 1 00 $q_{X3}$ $+$
Table 4: The short list of particle contents with an extended Higgs sector. $\phi_{1}^{\prime}$ and $\phi_{3}$ provide the masses of NG bosons to be heavy psuedo scalars. A $Z_{2}$ parity distinguishes $\phi_{1}$ and $\phi_{1}^{\prime}$, and let DM stable.
Appendix B The decay rate of $P$